OpenStudy (anonymous):

let f(x)=x+8 and g(x)=x^2-6x-7. find f(g(2))

OpenStudy (anonymous):

very similar to the last one again it takes two steps, not one

OpenStudy (anonymous):

first you need to find \(g(2)\)

OpenStudy (anonymous):

ok i did the problem and got 10 am i wrong because i could try again

OpenStudy (anonymous):

ok lets take it step by step \[g(x)=x^2-6x-7\] so \[g(2)=2^2-6\times 2-7\]

OpenStudy (anonymous):

when you compute that number you get \[2^2-6\times 2-7=4-12-7=-8-7=-15\] and that is just step one we still have to do step 2

OpenStudy (anonymous):

oh ok so what do we do for step 2?

OpenStudy (anonymous):

for step 2 we have to compute \[f(-15)\]

OpenStudy (anonymous):

\[f(x)=x+8\] so \[f(-15)=-15+8\]

OpenStudy (anonymous):

so its just 8 now

OpenStudy (anonymous):

because we minus the 15

OpenStudy (anonymous):

no what is \(-15+8\) ?

OpenStudy (anonymous):

-7

OpenStudy (anonymous):

right! and that is your final answer lets put it all together in one line

OpenStudy (anonymous):

\[f(g(2))=f(-15)=-7\]

OpenStudy (anonymous):

oh wow thanks i have another one if you could be so kind as to help me :)

OpenStudy (anonymous):

sure what is it?

OpenStudy (anonymous):

let f(x)=9x-2 and g(x)=-x+3. find f(g(x))

OpenStudy (anonymous):

ok for this one we have no numbers so we have to do something a bit different

OpenStudy (anonymous):

\[f(x)=9x-2,g(x)=-x+3\]

OpenStudy (anonymous):

\[f(g(x))\] always work from the inside out, it is easier in this case \(g(x)=-x+3\) so replace \(g(x)\) by \(-x+3\) to get \[f(g(x))=f(-x+3)\]

OpenStudy (anonymous):

then \[f(x)=9x-2\] or \[f(\heartsuit)=9\heartsuit-2\] we are going to replace \(\heartsuit\) by \((-x+3)\) to get \[f(\heartsuit)=f(-x+3)=9(-x+3)-2\]

OpenStudy (anonymous):

so is the answer -9x-2

OpenStudy (anonymous):

now we have to clean this up with a bit of algebra \[9(-x+3)-2=-9x+27-2=-9x+25\]

OpenStudy (anonymous):

don't forget the distributive law here

OpenStudy (anonymous):

\[9(-x+3)=9\times (-x)+9\times 3=-9x+27\]

OpenStudy (anonymous):

so it would be -9x+27

OpenStudy (anonymous):

then subtract the 2 to get a final answer of \(-9x+25\)

OpenStudy (anonymous):

oh ok i get that i have another one but its a little different and i need some help

OpenStudy (anonymous):

k

OpenStudy (anonymous):

let f(x)=-4x+7 and g(x)=10x-6 find f(g(x))

OpenStudy (anonymous):

no it is not really different \[f(g(x))=f(10x-6)=-4(10x-6)+7\] is the first two steps

OpenStudy (anonymous):

then some algebra to clean it up

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

you got the algebra?

OpenStudy (anonymous):

my algebra says the first part should be -40x

OpenStudy (anonymous):

but im not sure about the last part

OpenStudy (anonymous):

distribute the \(-4\) through the parentheses first to get \[-4(10x-6)+7=-40x+24-7\]

OpenStudy (anonymous):

then since \(24-7=17\) you end with \[-40x+17\]

OpenStudy (anonymous):

then wat do we do?

OpenStudy (anonymous):

wouldnt it be -40x+24

OpenStudy (anonymous):

no because we have that \(-7\) out at the end when you distribute, yes, you get \(-40x+24\) but there is also the \(-7\) giving a final answer of \[-40x+17\]

OpenStudy (anonymous):

ok thanks :)

OpenStudy (anonymous):

yw